1. Order: Amoxil 60mg PO q8h. The child weighs 20lbs and available dose is 125mg/5ml. The dose is 20-40mg/kg/day in 3 divided doses.

a. determined if the dose is safe.
b. if it is safe, determine how many ml would the nurse administer per dose.

2. Order: Marinol 5mg PO now. the patient is 10 years old that weighs 35kg and is 50 inches tall. the usual dose is 5mg/m². available dose is 2.5 capsule.

a. determine if dose is safe.
b. if it safe, how many capsules for dose.

need step by step calculations
Answer #1

ANSWER.

1) a, Calculation of safe dose for the patient.

             Ordered dose = 60 mg q 8h.

     q8h is 3 times a day.

       So

           Ordered dose = 60 * 3

                                  = 180 mg/day.

       Weight of the patient = 20 lbs.

   Convert lbs to kg for calculation

            1 lbs = 0.454 kg.

      Weight of the patient = 20 * 0.454

                                        = 9.08 kg.

     Recommended dose = 20-40 mg/kg/day.

Dose required for a patient weight 9.08 kg = 20 * 9.08 to 40 * 9.08

                                                                     = 181.6 mg to 363.2 mg.

* So ordered dose (180mg) is safe for this patient.

b) Calculation of volume(ml) of drug administer per dose.

           Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q).

     Here

            Desired dose(D) = 60 mg.

            Dose in hand(H) = 125 mg.

                Quantity(Q) = 5 ml.

As per above formula

      Volume(ml) of drug required = D/H * Q

                                                    = 60/125 * 5

                                                    = 0.48 * 5

                                                    = 2.4 ml per dose.

2) Calculation of safe dose.

         Ordered dose = 5 mg.

         Recommended dose = 5 mg/m2.

      Weight of the patient = 35 kg.

      Height of the patient = 50 inch.

Convert inch to cm for calculation

            1 inch = 2.5 cm.

   So

       Height of the patient = 50 * 2.5

                                        = 125 cm.

Body surface area (m2) of the patient

      Body surface area(BSA) = Squareroot (Height(cm) * Weight(kg))/3600)

                                              = Squareroot (35 * 125)/3600).

                                              = Squareroot (4375/3600)

                                              = Square root (1.3)

                                              = 1.14

                                              = 1 m2.

    Dose required for this patient = 5 * 1

                                                    = 5 mg.

* So ordered dose(5mg) is safe for this patient.

b) Calculation of number of capsules required.

            Ordered dose = 5 mg.

            Available dose = 2.5 mg/capsule.

   Number of capsules required = Ordered dose/Available dose.

                                                    = 5/2.5

                                                    = 2 capsules.